OpenStudy (anonymous):

for each pair , write the equivalent fractions that have the same denominator.....1/2 and 2/3

OpenStudy (anonymous):

3/6 and 4/6

OpenStudy (anonymous):

In that case if you want the same denominator the least common one is 6. 1/2 = 3/6 2/3 = 4/6

OpenStudy (anonymous):

thank you